# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v5

@@ 3 Problem Formulation
where $\{\varepsilon_{i}^{(t)}\}_{t\in\mathcal{T}}$ are (zero mean) $\sigma_{i}$-sub-gaussian demand noises following independent and identical distributions ($\varepsilon_{i}^{(t)}$ and $\varepsilon_{j}^{(t)}$ can be correlated with $i\neq j$, $i,j\in\mathcal{N}$), and $\bm{\theta}_{i}$ is an unknown vector of dimension $N$ with $i$-th entry equal to $\beta_{i}$ and the remaining entries being the values of $\bm{\gamma}_{i}\in\mathbb{R}^{N-1}$, ordered. The parameter vector $\bm{\gamma}_{i}$ measures how seller $i$ ’s demand is affected by competitor prices. We assume that the parameter space is such that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\bm{\theta}_{i},\psi_{i}\}_{i\in\mathcal{N}}$; a similar assumption is found in birge2024interfere; li2024lego. The above conditions hold if $\psi_{i},\psi_{i}^{\prime}>0$ and $\beta_{i}>0$, which are explicitly assumed in ˜3.1.
@@
The constants $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are known by the seller $i$ (they are used to compute the optimization problem in equation 11), however, firms can estimate them easily from historical sales data and operational capacity limits; moreover loose bounds suffice and do not affect the rates in Theorem˜5.4. The condition $\|\bm{\theta}_{i}\|_{2}=1$ guarantees the differentiability of the monotone single index model (chmielewski1981elliptically; balabdaoui2019least). Note that Equation˜6 is well defined, that is $\langle\bm{\theta}_{i},\mathbf{p}^{(t)}\rangle\in\mathcal{U}$, indeed $\|\bm{\theta}_{i}\|_{2}=1\implies|\langle\bm{\theta}_{i},\mathbf{p}\rangle|\leq\|\mathbf{p}\|_{2}\leq p_{\max}$, $\forall\mathbf{p}\in\mathcal{P}$. Regarding $\bm{\gamma}_{i}$, while many applications assume $\gamma_{ij}>0$, this restriction is not required for our algorithm or theoretical guarantees - our results remain valid when $\gamma_{ij}>0$. We allow $\gamma_{ij}$ to be negative to capture settings with negative cross-effects, such as vertically differentiated products, where raising a competitor’s price may reduce my demand if consumers perceive their product as higher quality.
@@
where $\bm{\Gamma}$ is called Best Response Operator, and its components $\Gamma_{i}$ are defined in equation 7.
@@ 3.1 Main Assumptions
The objective of this paper is to design an algorithm that guarantees sublinear regret for every seller, i.e., $\operatorname{Reg}_{i}(T)=o(T)$ for every $i\in\mathcal{N}$. Our policy consists of two phases: an exploration phase, in which each seller learns their individual best-response map $\hat{\Gamma}_{i}$ by consistently estimating the parameters $(\bm{\theta}_{i},\psi_{i})$, and an exploitation phase, in which, at each round $t$, seller $i$ sets their price according to $p_{i}^{(t)}=\hat{\Gamma}_{i}(\mathbf{p}_{-i}^{(t-1)})$, or equivalently $\mathbf{p}^{(t)}=\hat{\bm{\Gamma}}(\mathbf{p}^{(t-1)})$. The link between regret and equilibrium is straightforward: the individual regret is controlled if the iterates converge to the NE. In fact, $\operatorname{Reg}_{i}(T)\lesssim T\mathbb{E}\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}$, were $\mathbf{p}^{\star}$ is a NE (see equation 37 for a detailed bound). By the triangle inequality, we have
@@
where $L_{\bm{\Gamma}}$ is the Lipschitz constant of $\bm{\Gamma}$ and $\|\mathbf{F}\|_{\infty}=\sup_{\mathbf{p}\in\mathcal{P}}\|\mathbf{F}(\mathbf{p})\|$. Therefore, convergence to $\mathbf{p}^{\star}$ requires two ingredients: $(i)$ Contraction: $0<L_{\bm{\Gamma}}<1$ which ensures that deviations from the NE shrink geometrically; $(ii)$ Consistent estimation for $\|\hat{\bm{\Gamma}}-\bm{\Gamma}\|_{\infty}$, which follows from consistent estimation of $(\bm{\theta}_{i},\psi_{i})$ and strong concavity of $\operatorname{rev}_{i}(\cdot\mid\mathbf{p}_{-i})$ uniformly in $\mathbf{p}_{-i}$, for each seller $i\in\mathcal{N}$. Putting these together, we obtain (see equation G.2 for a formal inequality):
@@
By twice differentiability of $\psi_{i}$, ˜3.2 is guaranteed if $\mu_{i}\triangleq 2\beta_{i}\underline{B}_{\psi_{i}^{\prime}}-\beta_{i}^{2}\bar{p}_{i}B_{\psi_{i}^{\prime\prime}}>0$. Indeed, since by ˜3.1 $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ and $\psi_{i}^{\prime}\geq B_{\psi^{\prime}_{i}}$, we have
@@
˜3.2 guarantees that the best-response map $\Gamma_{i}$ in equation 7 is well defined for every $\mathbf{p}_{-i}$. More importantly, it guarantees the existence of the Nash equilibrium, in line with standard results in competitive games (scutari2014real; li2024lego; tsekrekos2024variational). Instances where ˜3.2 is satisfied include linear demand models (li2024lego), concave demand specifications, and a class of $s_{i}$-concave demand functions with $s_{i}>-1$. A more detailed discussion of these and further examples is provided in Section˜I.1. The next result is an immediate consequence of Theorem 3, in scutari2014real.
@@
Under assumptions equation 3.1 and equation 3.2, there exists a $\mathbf{p}^{*}\in\mathcal{P}$ satisfying the fixed point equation in equation 8.
By ˜3.2, the map $\Gamma_{i}(\mathbf{p}_{-i})$ can be recovered by solving the first order conditions (FOCs), $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i})=0$, and projecting the solution onto $\mathcal{P}_{i}$. We now determine a shape constraint assumption that is sufficient to satisfy the FOCs. Specifically, we can show that $\Gamma_{i}$ can be written as (see the proof of Lemma˜3.7 for the derivation)
@@
provided $\varphi_{i}^{-1}$ exists. Here $\Pi_{\mathcal{P}_{i}}$ is the projection into $\mathcal{P}_{i}$ and $\varphi_{i}$ is called virtual valuation function of firm $i$. For this purpose, we introduce ˜3.4, which makes the $\varphi_{i}$’s invertible.
@@
˜3.4 covers the linear demand model by li2024lego for which $\psi_{i}(u)=\alpha_{i}+u\implies\varphi_{i}(u)=2u+\alpha_{i}\implies\psi_{i}^{\prime}\geq 2$ and can also be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in fan2024policy, Assumption 1 in chen2018robust, Equation 1 in cole2014sample and golrezaei2019dynamic; javanmard2017perishability; javanmard2019dynamic which assume $\psi_{i}$ and $1-\psi_{i}$ to be log-concave (which specifically implies $c_{i}\geq 1$), where $\psi_{i}$ is a survival function.
We now present an equivalent formulation of ˜3.4 in terms of $s$-concavity, a condition that will allow us to estimate $\psi_{i}$ via shape constraints (the proof is provided in Appendix˜B). Examples of $s$-concave functions can be found in Appendix˜L.
@@
˜3.6 generalizes Assumption 2 in li2024lego (in their case $\psi_{i}(u)=u+\alpha_{i}\implies g_{i}(u)=(u+\alpha_{i})/2\implies g_{i}^{\prime}=1/2$, matching the contraction constant $L_{\bm{\Gamma}}$ in Equation 30 of their work). Similar assumptions are found in kachani2007modeling. The proof of the following result is relegated to Appendix˜C. In practice, ˜3.6 states that the influence of competitors’ prices on seller $i$’s optimal response is sufficiently small relative to the sensitivity to its own price.
@@
In this case ˜3.6 reduces to $\beta_{i}>\|\bm{\gamma}_{i}\|_{1}$, but since $\beta_{i}=(1-\|\bm{\gamma}_{i}\|_{2}^{2})^{1/2}$, the condition becomes $\|\bm{\gamma}_{i}\|_{1}^{2}+\|\bm{\gamma}_{i}\|_{2}^{2}<1$. But since $\|\bm{\gamma}_{i}\|_{2}\leq\|\bm{\gamma}_{i}\|_{1}$, a sufficient condition is $\|\bm{\gamma}_{i}\|_{1}<\nicefrac{{1}}{{\sqrt{2}}}$.
@@ 4 Proposed Algorithm
Our Algorithm˜1 consists of an initial exploration phase for parameter estimation, followed by an exploitation phase where sellers set prices based on the learned demand model. In the exploration phase, each seller $i\in\mathcal{N}$ samples prices $p_{i}^{(t)}$ from a distribution $\mathscr{D}_{i}$ for $t=1,2,\dots,\tau$. The common exploration length is $\tau\propto T^{\xi}$ for some $\xi\in(0,1)$, to be specified later (Remark˜5.6).
@@
See Section˜I.2 for a clear discussion of why sellers have no incentive to alter the exploration phase.
@@
Owing to space constraints, we defer to Section˜I.3.
@@
For every seller $i\in\mathcal{N}$, the parameter $\kappa_{i}$ can be chosen to minimize that seller’s total expected regret. By Theorem˜5.4 there is a unique optimum, denoted $\kappa_{i}^{\star}=\kappa_{i}^{\star}(N,\tau)\in(0,1)$ that is characterized by Equation˜14. Choosing $\kappa_{i}=\kappa_{i}^{\star}$ improves only the leading constant of the regret but leaves the convergence rate unchanged. For the exploration horizon $\tau$ we set $\tau\propto T^{\xi}$ for some $\xi>0$. Remark˜5.6 argues that $\xi^{\star}=5/7$ is the unique value that minimizes the total expected regret across all sellers. These values $\kappa_{i}^{*}$ and $\xi^{\star}$ are “optimal" only in the sense that they aim to minimize the regret constant and the exponent of $T$ respectively within our derived upper bound; we do not claim minimax optimality of the regret rate itself.
@@
    Exploration phase part 2: $\mathcal{T}_{i}^{(2)}$. Each seller $i$ defines $w^{(t)}_{i}\triangleq\langle\widetilde{\bm{\theta}}_{i},\textbf{p}^{(t)}\rangle$ for $t\in\mathcal{T}_{i}^{(2)}$ (note that $w_{i}^{(t)}\in\mathcal{U}$ because $\|\widetilde{\bm{\theta}}_{i}\|_{2}=1$, where $\mathcal{U}$ is defined in equation 5) and the estimator $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}=h_{s_{i}}\circ\widehat{\phi}_{i,\widetilde{\bm{\theta}}_{i}}$, where $s_{i}=c_{i}-1$ (known by ˜3.4), and
@@
    where $h_{s_{i}}$ is defined in equation 4. The class $\mathcal{H}_{i}$ consists of all functions $\phi:\mathcal{U}\to[\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ that are both monotone non-decreasing and concave. The bounds $\underline{B}_{\phi_{i}}$ and $\bar{B}_{\phi_{i}}$ are chosen so that $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}]=h_{s_{i}}([\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}])$, and are therefore known, indeed $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ (defined in ˜3.1) are known. A detailed justification of this estimator in equation 11 is provided in Section˜D.2.
@@
where $\mathbf{p}_{-i}^{(t-1)}$ are the prices of all the other firms at time t1t-1, that are made public. A price $p^{(t)}_{i}$ is then offered by firm $i$ as a maximizer of equation 12 over $\mathcal{P}_{i}$. We present the full procedure in Algorithm˜1 and a visual representation of the exploration-exploitation scheme in Figure˜9 in Appendix˜K. A summary of the information available to each seller to run Algorithm˜1 can be found in Appendix˜A.
@@
  Sample $\mathbf{p}^{(t)}=(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{N}^{(t)})\sim\mathscr{D}$, and let $y_{i}^{(t)}$ denote seller $i$’s demand as in equation 6.
@@
Each seller $i\in\mathcal{N}$ constructs an estimator $(\widetilde{\bm{\theta}}_{i},\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\cdot))$ using data $\{(\mathbf{p}^{(t)},y_{i}^{(t)})\}_{t=1,2,\ldots,\tau}$ as defined in equation 10 and equation 11
@@
  $(2)$ After all prices have been posted, each seller observes their demand $y_{i}^{(t)}$ as in equation 6.

### v6

@@ 3 Problem Formulation
where $\{\varepsilon_{i}^{(t)}\}_{t\in\mathcal{T}}$ are (zero mean) $\sigma_{i}$-sub-gaussian demand noises following independent and identical distributions ($\varepsilon_{i}^{(t)}$ and $\varepsilon_{j}^{(t)}$ can be correlated with $i\neq j$, $i,j\in\mathcal{N}$), and $\bm{\theta}_{i}$ is an unknown vector of dimension $N$ with $i$-th entry equal to $\beta_{i}$ and the remaining entries being the values of $\bm{\gamma}_{i}\in\mathbb{R}^{N-1}$, ordered. The parameter vector $\bm{\gamma}_{i}$ measures how seller $i$ ’s demand is affected by competitor prices. We assume that the parameter space is such that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\bm{\theta}_{i},\psi_{i}\}_{i\in\mathcal{N}}$; a similar assumption is found in Birge et al. (2024); Li et al. (2024). The above conditions hold if $\psi_{i},\psi_{i}^{\prime}>0$ and $\beta_{i}>0$, which are explicitly assumed in 3.1.
@@
The constants $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are known by the seller $i$ (they are used to compute the optimization problem in (11)), however, firms can estimate them easily from historical sales data and operational capacity limits; moreover loose bounds suffice and do not affect the rates in Theorem 5.4. The condition $\|\bm{\theta}_{i}\|_{2}=1$ guarantees the differentiability of the monotone single index model (Chmielewski, 1981; Balabdaoui et al., 2019). Note that Equation 6 is well defined, that is $\langle\bm{\theta}_{i},\mathbf{p}^{(t)}\rangle\in\mathcal{U}$, indeed $\|\bm{\theta}_{i}\|_{2}=1\implies|\langle\bm{\theta}_{i},\mathbf{p}\rangle|\leq\|\mathbf{p}\|_{2}\leq p_{\max}$, $\forall\mathbf{p}\in\mathcal{P}$. Regarding $\bm{\gamma}_{i}$, while many applications assume $\gamma_{ij}>0$, this restriction is not required for our algorithm or theoretical guarantees - our results remain valid when $\gamma_{ij}>0$. We allow $\gamma_{ij}$ to be negative to capture settings with negative cross-effects, such as vertically differentiated products, where raising a competitor’s price may reduce my demand if consumers perceive their product as higher quality.
@@
where $\bm{\Gamma}$ is called Best Response Operator, and its components $\Gamma_{i}$ are defined in (7).
@@ 3.1 Main Assumptions
The objective of this paper is to design an algorithm that guarantees sublinear regret for every seller, i.e., $\operatorname{Reg}_{i}(T)=o(T)$ for every $i\in\mathcal{N}$. Our policy consists of two phases: an exploration phase, in which each seller learns their individual best-response map $\hat{\Gamma}_{i}$ by consistently estimating the parameters $(\bm{\theta}_{i},\psi_{i})$, and an exploitation phase, in which, at each round $t$, seller $i$ sets their price according to $p_{i}^{(t)}=\hat{\Gamma}_{i}(\mathbf{p}_{-i}^{(t-1)})$, or equivalently $\mathbf{p}^{(t)}=\hat{\bm{\Gamma}}(\mathbf{p}^{(t-1)})$. The link between regret and equilibrium is straightforward: the individual regret is controlled if the iterates converge to the NE. In fact, $\operatorname{Reg}_{i}(T)\lesssim T\mathbb{E}\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}$, were $\mathbf{p}^{\star}$ is a NE (see (37) for a detailed bound). By the triangle inequality, we have
@@
where $L_{\bm{\Gamma}}$ is the Lipschitz constant of $\bm{\Gamma}$ and $\|\mathbf{F}\|_{\infty}=\sup_{\mathbf{p}\in\mathcal{P}}\|\mathbf{F}(\mathbf{p})\|$. Therefore, convergence to $\mathbf{p}^{\star}$ requires two ingredients: $(i)$ Contraction: $0<L_{\bm{\Gamma}}<1$ which ensures that deviations from the NE shrink geometrically; $(ii)$ Consistent estimation for $\|\hat{\bm{\Gamma}}-\bm{\Gamma}\|_{\infty}$, which follows from consistent estimation of $(\bm{\theta}_{i},\psi_{i})$ and strong concavity of $\operatorname{rev}_{i}(\cdot\mid\mathbf{p}_{-i})$ uniformly in $\mathbf{p}_{-i}$, for each seller $i\in\mathcal{N}$. Putting these together, we obtain (see (G.2) for a formal inequality):
@@
By twice differentiability of $\psi_{i}$, 3.2 is guaranteed if $\mu_{i}\triangleq 2\beta_{i}\underline{B}_{\psi_{i}^{\prime}}-\beta_{i}^{2}\bar{p}_{i}B_{\psi_{i}^{\prime\prime}}>0$. Indeed, since by 3.1 $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ and $\psi_{i}^{\prime}\geq B_{\psi^{\prime}_{i}}$, we have
@@
3.2 guarantees that the best-response map $\Gamma_{i}$ in (7) is well defined for every $\mathbf{p}_{-i}$. More importantly, it guarantees the existence of the Nash equilibrium, in line with standard results in competitive games (Scutari et al., 2014; Li et al., 2024; Tsekrekos and Yannacopoulos, 2024). Instances where 3.2 is satisfied include linear demand models (Li et al., 2024), concave demand specifications, and a class of $s_{i}$-concave demand functions with $s_{i}>-1$. A more detailed discussion of these and further examples is provided in Section I.1. The next result is an immediate consequence of Theorem 3, in Scutari et al. (2014).
@@
Under Assumptions (3.1) and (3.2), there exists a $\mathbf{p}^{*}\in\mathcal{P}$ satisfying the fixed point equation in (8).
By 3.2, the map $\Gamma_{i}(\mathbf{p}_{-i})$ can be recovered by solving the first order conditions (FOCs), $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i})=0$, and projecting the solution onto $\mathcal{P}_{i}$. We now determine a shape constraint assumption that is sufficient to satisfy the FOCs. Specifically, we can show that $\Gamma_{i}$ can be written as (see the proof of Lemma 3.7 for the derivation)
@@
provided $\varphi_{i}^{-1}$ exists. Here $\Pi_{\mathcal{P}_{i}}$ is the projection into $\mathcal{P}_{i}$ and $\varphi_{i}$ is called virtual valuation function of firm $i$. For this purpose, we introduce 3.4, which makes the $\varphi_{i}$’s invertible.
@@
3.4 covers the linear demand model by Li et al. (2024) for which $\psi_{i}(u)=\alpha_{i}+u\implies\varphi_{i}(u)=2u+\alpha_{i}\implies\psi_{i}^{\prime}\geq 2$ and can also be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in Fan et al. (2024), Assumption 1 in Chen and Farias (2018), Equation 1 in Cole and Roughgarden (2014) and Golrezaei et al. (2019); Javanmard (2017); Javanmard and Nazerzadeh (2019) which assume $\psi_{i}$ and $1-\psi_{i}$ to be log-concave (which specifically implies $c_{i}\geq 1$), where $\psi_{i}$ is a survival function.
We now present an equivalent formulation of 3.4 in terms of $s$-concavity, a condition that will allow us to estimate $\psi_{i}$ via shape constraints (the proof is provided in Appendix B). Examples of $s$-concave functions can be found in Appendix L.
@@
3.6 generalizes Assumption 2 in Li et al. (2024) (in their case $\psi_{i}(u)=u+\alpha_{i}\implies g_{i}(u)=(u+\alpha_{i})/2\implies g_{i}^{\prime}=1/2$, matching the contraction constant $L_{\bm{\Gamma}}$ in Equation 30 of their work). Similar assumptions are found in Kachani et al. (2007). The proof of the following result is relegated to Appendix C. In practice, 3.6 states that the influence of competitors’ prices on seller $i$’s optimal response is sufficiently small relative to the sensitivity to its own price.
@@
In this case 3.6 reduces to $\beta_{i}>\|\bm{\gamma}_{i}\|_{1}$, but since $\beta_{i}=(1-\|\bm{\gamma}_{i}\|_{2}^{2})^{1/2}$, the condition becomes $\|\bm{\gamma}_{i}\|_{1}^{2}+\|\bm{\gamma}_{i}\|_{2}^{2}<1$. But since $\|\bm{\gamma}_{i}\|_{2}\leq\|\bm{\gamma}_{i}\|_{1}$, a sufficient condition is $\|\bm{\gamma}_{i}\|_{1}<\nicefrac{{1}}{{\sqrt{2}}}$.
@@ 4 Proposed Algorithm
Our Algorithm 1 consists of an initial exploration phase for parameter estimation, followed by an exploitation phase where sellers set prices based on the learned demand model. In the exploration phase, each seller $i\in\mathcal{N}$ samples prices $p_{i}^{(t)}$ from a distribution $\mathscr{D}_{i}$ for $t=1,2,\dots,\tau$. The common exploration length is $\tau\propto T^{\xi}$ for some $\xi\in(0,1)$, to be specified later (Remark 5.6).
@@
See Section I.2 for a clear discussion of why sellers have no incentive to alter the exploration phase.
@@
Owing to space constraints, we defer to Section I.3.
@@
For every seller $i\in\mathcal{N}$, the parameter $\kappa_{i}$ can be chosen to minimize that seller’s total expected regret. By Theorem 5.4 there is a unique optimum, denoted $\kappa_{i}^{\star}=\kappa_{i}^{\star}(N,\tau)\in(0,1)$ that is characterized by Equation 14. Choosing $\kappa_{i}=\kappa_{i}^{\star}$ improves only the leading constant of the regret but leaves the convergence rate unchanged. For the exploration horizon $\tau$ we set $\tau\propto T^{\xi}$ for some $\xi>0$. Remark 5.6 argues that $\xi^{\star}=5/7$ is the unique value that minimizes the total expected regret across all sellers. These values $\kappa_{i}^{*}$ and $\xi^{\star}$ are “optimal” only in the sense that they aim to minimize the regret constant and the exponent of $T$ respectively within our derived upper bound; we do not claim minimax optimality of the regret rate itself.
@@
    Exploration phase part 2: $\mathcal{T}_{i}^{(2)}$. Each seller $i$ defines $w^{(t)}_{i}\triangleq\langle\widetilde{\bm{\theta}}_{i},\textbf{p}^{(t)}\rangle$ for $t\in\mathcal{T}_{i}^{(2)}$ (note that $w_{i}^{(t)}\in\mathcal{U}$ because $\|\widetilde{\bm{\theta}}_{i}\|_{2}=1$, where $\mathcal{U}$ is defined in (5)) and the estimator $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}=h_{s_{i}}\circ\widehat{\phi}_{i,\widetilde{\bm{\theta}}_{i}}$, where $s_{i}=c_{i}-1$ (known by 3.4), and
@@
    where $h_{s_{i}}$ is defined in (4). The class $\mathcal{H}_{i}$ consists of all functions $\phi:\mathcal{U}\to[\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ that are both monotone non-decreasing and concave. The bounds $\underline{B}_{\phi_{i}}$ and $\bar{B}_{\phi_{i}}$ are chosen so that $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}]=h_{s_{i}}([\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}])$, and are therefore known, indeed $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ (defined in 3.1) are known. A detailed justification of this estimator in (11) is provided in Section D.2.
@@
where $\mathbf{p}_{-i}^{(t-1)}$ are the prices of all the other firms at time t1t-1, that are made public. A price $p^{(t)}_{i}$ is then offered by firm $i$ as a maximizer of (12) over $\mathcal{P}_{i}$. We present the full procedure in Algorithm 1 and a visual representation of the exploration-exploitation scheme in Figure 9 in Appendix K. A summary of the information available to each seller to run Algorithm 1 can be found in Appendix A.
@@
  Sample $\mathbf{p}^{(t)}=(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{N}^{(t)})\sim\mathscr{D}$, and let $y_{i}^{(t)}$ denote seller $i$’s demand as in (6).
@@
Each seller $i\in\mathcal{N}$ constructs an estimator $(\widetilde{\bm{\theta}}_{i},\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\cdot))$ using data $\{(\mathbf{p}^{(t)},y_{i}^{(t)})\}_{t=1,2,\ldots,\tau}$ as defined in (10) and (11)
@@
  $(2)$ After all prices have been posted, each seller observes their demand $y_{i}^{(t)}$ as in (6).